# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v2

@@ 4.1 Tasks and datasets
We utilize two established benchmarks Cogenesis (Zhang et al., 2024b) and Longlamp (Kumar et al., 2024). Cogenesis provides summarized user profiles and past experiences, with the aim of generating highly personalized content. We evaluate on its official test set. Longlamp focuses on personalized long-form generation, where each writing query includes ground truth responses and the user’s previous writing records. We augment each instance with the top-5 relevant historical records retrieved by bge-reranker-v2-m3 (Chen et al., 2024) as its personal context. We conduct experiments on the official test sets of three constituent tasks: abstract generation (Tang et al., 2008), review writing (Ni et al., 2019), and topic writing (Völske et al., 2017).
@@
We evaluate on four datasets: Helpsteer (Wang et al., 2023b), Truthful QA (Lin et al., 2022), UltraChat (Ding et al., 2023), and Personal Preference Eval (Gao et al., 2024). These datasets require generating content aligned with explicitly stated user preferences. Due to the large scale of these datasets, we randomly sample 200 test instances from each dataset.
Detailed dataset description and examples can be found in Appendix B.5
@@ 4.2 Evaluation metrics
To ensure fairness in comparison, we employ task-specific metrics proposed by these datasets themselves for evaluation. For Cogenesis, we use GPT-4o-2024-08-06 to evaluate the response’s overall and personalized scores, averaging the results over five runs with temperature set to 0 to mitigate potential instability. For Longlamp, we employ ROUGE (Lin, 2004) and METEOR (Banerjee & Lavie, 2005) scores to measure content overlap and semantic alignment, supplementing them with human evaluation (detailed in Appendix B.6) to account for the potential limitations of these automatic metrics. For preference alignment tasks, following prior work (Zhang et al., 2025b; Zhong et al., 2024), we set user preferences to be concise, creative, uplifting, and verbose, and utilize the reward model ArmoRM-8B (Wang et al., 2024) to evaluate the degree to which the generations align with these preferences.
4.3 Settings and baselines
To verify the robustness of our framework, we evaluate five model pairs of varying scales and architectures, including (1) Qwen2.5-7B-Instruct with Qwen2.5-1.5B-Instruct (2) Qwen2.5-32B-Instruct with Qwen2.5-7B-Instruct (Team, 2024) (3) Llama3.1-8B-Instruct with Llama-3.2-1B-Instruct (et al., 2024) (4) Qwen3-0.6B with Qwen3-8B and (5) Qwen3-0.6B with Qwen3-32B (Yang et al., 2025). Detailed parameters and settings are presented in Appendix B.1. We establish the following critical baselines per pair:
  • SLM w/o: Base performance of standalone SLMs without personal context.
  • SLM w/: SLMs augmented with personal context.
  • LLM w/o: Cloud-based LLMs without access to personal context.
  • LLM w/: The near-upper-bound performance where LLMs directly access personal context. It is important to note that this approach can lead to privacy breaches and does not align with our task setting.
@@ 4.3 Settings and baselines
Table 1: Comparative performance across eight personalized content generation and preference alignment tasks. Metrics include overall (Ovl) and personalized (Per) scores for Cogenesis, ROUGE-1/-L (R-1/-L) and METEOR (MET) for Longlamp datasets, and averaged alignment scores for four user preferences. Bold entries indicate that CoSteer outperforms the three baseline methods compared against. Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
@@ 4.4 Main results
Table 1 presents our core experimental findings. For clarity, here we present results from first three of our five model pair configurations and report the average performance across the four preference alignment tasks. The complete results, including for the remaining two model pairs (Table 6) and the detailed per-dataset preference alignment scores (Table 5), can be found in the Appendix B.2. Furthermore, we conducted paired t-tests to verify that our improvements are statistically significant, with a detailed analysis available in Appendix B.3.
Overall Performance
We conducted extensive experiments across two major task categories (personalized content generation and preference alignment), different model architectures (Llama and Qwen), and various model sizes (from 0.6B to 32B, see Table 1 and  6 ) . In the vast majority of settings, our method significantly outperforms key baselines: (1) cloud-based LLMs responding without access to personal context, and (2) local SLMs with or without such context. Remarkably, our performance often approaches or even exceeds the “near-upper-bound” achieved by LLMs with full context access. These results show that CoSteer enables cloud-based LLMs to generate personalized content using only locally stored user context. We further analyze the impact of task complexity and model scale in Appendix B.4, confirming the robustness and adaptability of our approach.
@@ 4.5 Comparison with alternative methods
  1. 1.
    Linear Alignment (Gao et al., 2024) and Context Steering(He et al., 2025) : These two methods are formally equivalent to scaling the logit difference of the SLM (with and without personal context) and applying it to its own inference process. This constitutes an inference-time optimization performed directly on the SLM.
  2. 2.
    Supervised Fine-Tuning (SFT): This involves directly fine-tuning the localized SLMs using personal data. Although we argue that fine-tuning a separate model for each user and task is impractical in real-world scenarios due to privacy and data constraints, we report its performance here for a thorough academic comparison.
  3. 3.
    Proxy-tuning Liu et al. (2024): This method leverages the difference in the SLM’s logit distributions before and after the SFT process described above. This difference is directly added to the LLM’s logits to steer its output distribution.
We conduct experiments on Longlamp using the Qwen7B-1.5B configuration. Implementation details and parameter settings for these baselines are provided in the Appendix B.8. As shown in Table 2, the results again confirm the superiority of the CoSteer framework. Our method adeptly combines the performance advantage of localized SLMs in handling private data with the powerful general capabilities of cloud-based LLMs, achieving the best performance among all compared methods without requiring any training.
@@ 4.6 Ablation Study and Parameter Analysis
Ablation study
We conduct an ablation study to validate the contributions of the key components in our framework. The SLM delta steering signal (the $\beta$ term in Eq. (4)) is the central mechanism of CoSteer. Our study therefore focuses on the impact of individually removing two other components: the LLM Alignment Term (the $\alpha$ term) and the KL Divergence Regularizer from the FTRL algorithm. The results in Table 2 proved their effectiveness. A broader ablation on the entire iterative FTRL (we name it LightCosteer) is discussed in Section 5.3.
@@
We examine the impact of different hyperparameters on our method. We conduct experiments on abstract generation using the Qwen7B-1.5B configuration, and evaluate with the ROUGE-L metric. When studying one parameter, we keep all others at their default values. Results and analysis presented in Appendix B.9 and Figure 2 show consistently strong performance across a broad range of values, confirming that the default hyperparameter settings (provided in Appendix B.1) are both robust and well-calibrated.
5 Discussion: Practical Deployment Considerations
In Section 4 , we showed the empirical superiority of the CoSteer framework through comprehensive experiments and comparisons. In this section, we shift our focus to its practical viability, analyzing its feasibility and performance from various angles that simulate real-world conditions and constraints.
@@ 5.2 Generalization across Model Scales and Architectures
A key advantage of our framework is its ability to generalize across diverse model configurations, a critical feature for real-world deployment. We validate this generalization capability in two dimensions: model scale and architecture.
@@
While pairing models from the same family is common, a practical framework should accommodate scenarios where the LLM and SLM differ in architecture and tokenizer. To this end, we investigated two strategies. The first, CoSteer_map, aligns models via vocabulary intersection, while effective, its applicability is limited when tokenizers are highly divergent. Therefore, we propose CoSteer_byte which operates on a shared byte representation, making it vocabulary-agnostic (Hayase et al., 2025). As shown in Table 11, both strategies successfully enable the SLM to guide the LLM, confirming that CoSteer is not confined to homogeneous model families. This flexibility demonstrates the framework’s generalization across architectures. Full implementation details are provided in Appendix 11.
@@ 5.3 Efficiency and Practical Variants
Finally, we analyze the efficiency and overhead of our framework. At a high level, a key advantage of CoSteer is its computational efficiency. Compared to methods that require concatenating long personal context to the LLM’s input (i.e., LLM w/), our framework offers significant computational savings. By processing the private context locally with an SLM, CoSteer substantially reduces the remote LLM’s workload. Our FLOPs estimation (Narayanan et al., 2021) in Table 14 shows that with a 1000-token context, the privacy-violating LLM w/ approach is approximately 3.3x more computationally expensive than CoSteer.
Although our closed-form solution is efficient, the full CoSteer framework still employs iterative optimization (e.g., $T=20$), which can introduce latency. Furthermore, each generation step requires communication between the local device and the server. To address these practical overheads, we propose and evaluate two streamlined variants:
  • LightCoSteer: To minimize computational overhead, this variant eliminates the iterative process by setting the number of optimization iterations $T=1$. This simplifies the framework to a single-step logit adjustment, providing a lightweight yet effective alternative.
  • AdaCoSteer: To reduce communication overhead, this variant adaptively deactivates steering. We observe that as generation proceeds, SLM and LLM predictions often converge. AdaCoSteer leverages this by terminating the steering process once the LLM’s token confidence exceeds a set threshold for a few consecutive steps (Song et al., 2025), allowing the LLM to complete the generation on its own.
@@ 6 Conclusion
In this work, we introduced CoSteer, a collaborative framework that enables real-time, privacy-preserving personalization for LLMs via localized delta steering. Our approach leverages an on-device SLM to guide a cloud-based LLM, effectively incorporating user context without transmitting sensitive data or requiring fine-tuning. Extensive experiments demonstrate that CoSteer offers a practical and effective solution for harmonizing the powerful general capabilities of LLMs with user-specific context and privacy.
Appendices Contents
Appendix A Methodology Detail
A.1 Related work
As shown in 3, CoSteer is unique as it is both collaborative and completely tuning-free. By formulating the steering as a local online learning problem, it allows for a lightweight, highly adaptive, and privacy protection framework.
@@ A.1 Related work
Method
Training Free? Weak-to-Strong Collaborative? No Additional Modules?
Linear Alignment (Gao et al., 2024) Yes No Yes
Contrastive Decoding (Li et al., 2023)
Context Steering (He et al., 2025)
Amulet (Zhang et al., 2025b)
PAD (Chen et al., 2025) Noa No No
Drift (Kim et al., 2025)
Proxy-tuning (Liu et al., 2024) Nob Yes Yes
Cogenesis (Zhang et al., 2024b) Noc Yes No
@@
A.2 Framework
Our CoSteer framework workflow is shown in Algorithm 1
Algorithm 1 CoSteer Framework
1:
2:Cloud-based LLM policy generator $\pi_{\text{LLM}}$, local SLM policy generator $\pi_{\text{SLM}}$, user query $p_{\text{base}}$, personal context $p_{\text{pers}}$, current sequence $s$, max tokens $M$, hyperparameters: T, $\alpha,\beta,\lambda,\eta>0$
3:Personalized sequence $s$
4:Initialize $s\leftarrow\emptyset$
5:repeat
6:  $\pi_{\text{base}}\leftarrow\pi_{\text{LLM}}(a|p_{\text{base}},s)$ $\triangleright$ Cloud computation
7:  $\pi^{*}_{\text{base}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},s)$ $\triangleright$ Edge computation
8:  $\pi^{*}_{\text{pers}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},p_{\text{pers}},s)$
9:  $\Delta\leftarrow\log\pi^{*}_{\text{pers}}-\log\pi^{*}_{\text{base}}$
10:  $\pi_{0}\leftarrow\pi_{\text{base}}$ $\triangleright$ Initialize policy
11:  for $t=1$ to $T$ do
12:   $u_{t}\leftarrow\alpha(\log\pi_{t-1}-\log\pi_{\text{base}})+\beta\Delta$
13:   Update policy using Equation 7
14:  end for
15:  $\pi^{*}\leftarrow\pi_{T}$ $\triangleright$ Final optimized policy
16:  Sample $a\sim\pi^{*}$, Update $s\leftarrow s\circ a$
17:until $\text{len}(s)\geq M$ or EOS generated
18:return $s$
A.3 Proof
Similar to Zhang et al. (2025b), We try to solve the closed-form solution of Equation 6:

### v3

@@ 4.1 Tasks and datasets
We utilize two established benchmarks, Cogenesis (cogenesis) and LongLaMP (longlamp). Cogenesis provides summarized user profiles and past experiences, with the aim of generating highly personalized content. We evaluate on its official test set. LongLaMP focuses on personalized long-form generation, where each writing query includes ground truth responses and the user’s previous writing records. We augment each instance with the top-5 relevant historical records retrieved by bge-reranker-v2-m3 (bgem3) as its personal context. We conduct experiments on the official test sets of three constituent tasks: abstract generation (abstract), review writing (review), and topic writing (topicwriting).
@@
We evaluate on four datasets: HelpSteer (helpsteer), Truthful QA (truthfulqa), UltraChat (ultrachat), and Personal Preference Eval (gao2024linearalignmentclosedformsolution). These datasets require generating content aligned with explicitly stated user preferences. Due to the large scale of these datasets, we randomly sample 200 test instances from each dataset.
Dataset description and examples are in Appendix B.5
@@ 4.2 Evaluation metrics
To ensure fairness in comparison, we employ task-specific metrics proposed by these datasets themselves for evaluation. For Cogenesis, we use GPT-4o-2024-08-06 to evaluate the response’s overall and personalized scores, averaging the results over five runs with temperature set to 0 to mitigate potential instability. For LongLaMP, we employ ROUGE (lin2004rouge) and METEOR (banerjee2005meteor) scores to measure content overlap and semantic alignment, supplementing them with human evaluation (detailed in Appendix B.6) to account for the potential limitations of these automatic metrics. For preference alignment tasks, following prior work (zhang2025amulet; zhong2024panaceaparetoalignmentpreference), we set user preferences to be concise, creative, uplifting, and verbose, and utilize the reward model ArmoRM-8B (armo) to evaluate the degree to which the generations align with these preferences.
@@
Qwen 8B-0.6B SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
Qwen 32B-0.6B SLM w/o 6.84 6.64 37.13 20.41 21.00 24.41 12.83 12.39 24.19 11.88 15.74 .5710 .5359 .5981 .5937
CoSteer 8.52 8.56 43.74 23.69 30.97 30.89 14.14 18.87 26.75 13.17 21.32 .7858 .7883 .6873 .7790
Table 1: Comparative performance across eight personalized content generation and preference alignment tasks. Metrics include overall (Ovl) and personalized (Per) scores for Cogenesis, ROUGE-1/-L (R-1/-L) and METEOR (MET) for Longlamp datasets, and averaged alignment scores for four user preferences. Bold entries indicate that CoSteer outperforms the three baseline methods. Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
4.3 Settings and baselines
To verify the robustness of our framework, we evaluate five model pairs of varying scales and architectures, including (1) Qwen2.5-7B-Instruct with Qwen2.5-1.5B-Instruct (2) Qwen2.5 -32B-Instruct with Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport) (3) Llama3.1-8B-Instruct with Llama3.2-1B-Instruct (grattafiori2024llama) (4) Qwen3-8B with Qwen3-0.6B and (5) Qwen3-32B with Qwen3-0.6B (qwen3). Detailed parameters and settings are presented in Appendix B.1.
We establish the following critical baselines per pair: (1) SLM w/o: Base performance of standalone SLMs without personal context. (2) SLM w/: SLMs augmented with personal context. (3) LLM w/o: Cloud-based LLMs without access to personal context. (4) LLM w/: The near-upper-bound performance where LLMs directly access personal context. It is important to note that this approach can lead to privacy breaches and does not align with our task setting.
@@ 4.4 Main Results
Table 1 summarizes our main results across all five model-pair configurations, reporting the average performance over four preference alignment tasks. Overall, CoSteer consistently improves preference alignment compared with both cloud-based LLMs without personal context and local SLMs with or without context, and these gains are statistically significant under paired t-tests. Detailed per-dataset scores are deferred to Appendix B.2 (Table 5), and the full significance analysis is provided in Appendix B.3.
Consistent and substantial gains. Across the vast majority of settings, CoSteer significantly outperforms two key baseline families: (1) cloud-based LLMs responding without personal context, and (2) local SLMs with or without context. This indicates that CoSteer can reliably translate locally stored user context into measurable improvements in personalized generation and preference alignment.
Robustness across standard model pairs. On common pairings (e.g., Qwen2.5 32B–7B and 7B–1.5B), the improvements are especially consistent and robust, suggesting that CoSteer is insensitive to a specific architecture or capacity range, and generalizes well under typical deployment scales.
Nuances driven by length and model-specific style. Llama 3 performance correlates with output length: it delivers strong gains on concise Preference Alignment tasks, while improvements are more moderate in long-form generation. We attribute this to the compact Llama-1B’s stylistic bias toward brevity, limiting expressiveness when the target requires sustained long-form writing.
Extreme capacity gaps and conservative steering. With the ultra-compact Qwen3-0.6B, CoSteer yields clear benefits on complex personalized writing, whereas gains on simpler alignment tasks are less pronounced. We hypothesize this stems from CoSteer’s regularization mechanism, which conservatively limits steering intensity under extreme capacity gaps, thereby preventing the tiny SLM from overly constraining the LLM and degrading global coherence.
Approaching the near-upper-bound. Remarkably, CoSteer often approaches or even exceeds the “near-upper-bound” achieved by LLMs with full context access. These results demonstrate that CoSteer enables cloud-based LLMs to generate personalized content using only locally stored user context. We further analyze the interactions among task complexity and model scale in Appendix B.4.
@@ 4.5 Comparison with Alternative Methods
(1) Linear Alignment (gao2024linearalignmentclosedformsolution) and Context Steering(he2025contextsteeringcontrollablepersonalization) : These two methods are formally equivalent to scaling the logit difference of the SLM (with and without personal context) and applying it to its own inference process. This constitutes an inference-time optimization performed directly on the SLM;
(2) Supervised Fine-Tuning (SFT): This involves directly fine-tuning the localized SLMs using personal data. Although we argue that fine-tuning a separate model for each user and task is impractical in real-world scenarios due to privacy and data constraints, we report its performance here for a thorough academic comparison;
(3) Proxy-tuning (liu2024tuninglanguagemodelsproxy): This method leverages the difference in the SLM’s logit distributions before and after the SFT process described above. This difference is added to the LLM’s logits to steer its output distribution.
We conduct experiments on LongLaMP using the Qwen7B–1.5B configuration (Table 2). For completeness, results for all five LLM–SLM pairs are reported in Appendix B.8 (Table 8). As quantitatively evidenced in Table 2, CoSteer demonstrates superior performance across the vast majority of experimental settings, and this competitive advantage remains robust across a wide spectrum of model pairs in Table 8. Notably, CoSteer combines the privacy-preserving strength of localized SLMs with the general capability of cloud LLMs, without any training.
@@ 4.6 Ablation Study and Parameter Analysis
Ablation studies
We conduct an ablation studies to validate the contributions of the key components in our framework. The SLM delta steering signal (the $\beta$ term in Eq. (4)) is the central mechanism of CoSteer. Our study therefore focuses on the impact of individually removing two other components: the LLM Alignment Term (the $\alpha$ term) and the KL Divergence Regularizer from the FTRL algorithm. The results in Table 2 proved their effectiveness. A broader ablation on the entire iterative FTRL (we name it LightCosteer) is discussed in Section 5.3.
@@
We examine the impact of different hyperparameters on our method. We conduct experiments on abstract generation using the Qwen7B-1.5B configuration, and evaluate with the ROUGE-L metric. When studying one parameter, we keep all others at default. Results and analysis presented in Appendix B.9 and Figure 2 show consistently strong performance across a broad range of values, confirming that the default hyperparameter settings (provided in Appendix B.1) are both robust and well-calibrated.
5 Practical Considerations for Deployment
In Section 4, we showed the empirical superiority of the CoSteer framework through extensive experiments and comparisons. In this section, we focus on its practical viability, analyzing its feasibility and performance from various angles that simulate real-world conditions and constraints.
@@ 5.2 Generalization across Model Scales and Architectures
Our framework generalizes to different model configurations, which is essential for practical deployment. We validate this in two dimensions: model scale and architecture.
@@
While pairing models from the same family is common, a practical framework should accommodate scenarios where the LLM and SLM differ in architectures and tokenizers. To this end, we investigated two strategies. The first, CoSteer_map, aligns models via vocabulary intersection, while effective, its applicability is limited when tokenizers are highly divergent. Therefore, we propose CoSteer_byte which operates on a shared byte representation, making it vocabulary-agnostic (hayase2025byte-sampling). As shown in Table 11, both strategies successfully enable the SLM to guide the LLM, confirming that CoSteer is not confined to homogeneous model families. This flexibility demonstrates the framework’s generalization across architectures. Full implementation details are provided in Appendix 11.
@@ 5.3 Efficiency and Practical Variants
Finally, we analyze the efficiency and overhead of our framework. At a high level, a key advantage of CoSteer is its computational efficiency. Compared to methods that require concatenating long personal context to the LLM’s input (i.e., LLM w/), our framework offers significant computational savings. By processing the private context locally with an SLM, CoSteer substantially reduces the remote LLM’s workload. Our FLOPs estimation (flop) in Table 14 shows that with a 1000-token context, the privacy-violating LLM w/ approach is approximately 3.3x more computationally expensive than CoSteer.
Although our closed-form solution is efficient, the full CoSteer framework still employs iterative optimization (e.g., $T=20$), which can introduce latency. Furthermore, each generation step requires communication between the local device and the server. As detailed in Figure 3 (Appendix), CoSteer uses an asynchronous two-stream pipeline where local SLM inference is largely masked by the network/cloud path; thus, the unmasked critical-path overhead is dominated by network transmission and the (iterative) local FTRL update. To address these practical overheads, we propose and evaluate two streamlined variants:
(1) LightCoSteer: To minimize computational overhead, this variant eliminates the iterative process by setting the number of optimization iterations $T=1$. This simplifies the framework to a single-step logit adjustment, providing a lightweight yet effective alternative.
(2) AdaCoSteer: To reduce communication overhead, this variant adaptively deactivates steering. We observe that as generation proceeds, SLM and LLM predictions often converge. AdaCoSteer leverages this by terminating the steering process once the LLM’s token confidence exceeds a threshold for a few consecutive steps (song-etal-2025-well-begun), allowing the LLM to complete the generation on its own.
@@ 6 Conclusion
In this work, we propose CoSteer, a collaborative frame- work that addresses the challenge of achieving real-time, privacy-preserving personalized generation by combining the strengths of local SLMs and cloud-based LLMs. Specif- ically, CoSteer formulates decoding-time personalization as efficient logit-level steering, using the on-device local logit delta between context-aware and context-agnostic SLM pre- dictions with a closed-form update rule. We further present streamlined variants (LightCoSteer and AdaCoSteer) to trade off computation and communication overhead for prac- tical edge deployment. We demonstrate the effectiveness and practical applicability of this framework through exten- sive experiments, showing its ability to deliver personalized content while maintaining privacy. Future work will inves- tigate broader applicability across domains and modalities, and further optimize scalability, latency, and resource usage for edge deployment.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Methodology Details
A.1 Related Work and Distinction
As summarized in Table 3, CoSteer occupies a unique technical position by being both collaborative and completely tuning-free. While it shares conceptual roots with inference-time steering methods, it introduces critical innovations tailored to the privacy-constrained cloud-edge scenario.
Distinction from Context Steering Methods
Unlike methods like Context Steering (he2025contextsteeringcontrollablepersonalization) or Linear Alignment (gao2024linearalignmentclosedformsolution), which operate in a single-model setting where a model steers itself, CoSteer addresses a fundamentally different challenge: heterogeneous cross-model collaboration. We demonstrate a novel “weak-to-strong” capability where a tiny, local SLM effectively steers a massive, remote LLM. This allows high-quality personalization without exposing private data to the cloud, a constraint that single-model cloud deployment cannot satisfy.
Technical Novelty beyond Optimization
While we adopt the FTRL algorithm as our solver, our core contribution lies in the problem formulation. We novelly define the utility function (Equation 4) based on local logit deltas to repurpose online learning for solving a distributed coordination problem. Furthermore, to address practical bottlenecks unique to this collaborative architecture, we introduce system-level innovations not found in prior work: AdaCoSteer minimizes communication latency via confidence-based termination, and Byte-Level Fusion resolves tokenizer mismatches to enable collaboration between heterogeneous model families (e.g., Llama and Qwen).
@@ A.1 Related Work and Distinction
Method Training Free? Weak-to-Strong Collaborative? No Additional Modules?
Linear Alignment (gao2024linearalignmentclosedformsolution) Yes No Yes
Contrastive Decoding (li2023contrastivedecodingopenendedtext)
Context Steering (he2025contextsteeringcontrollablepersonalization)
Amulet (zhang2025amulet)
PAD (pad) Noa No No
Drift (Kim2025DriftDP)
Proxy-tuning (liu2024tuninglanguagemodelsproxy) Nob Yes Yes
Cogenesis (cogenesis) Noc Yes No
@@
A.2 Proof
Similar to zhang2025amulet, We try to solve the closed-form solution of Equation 6: